OpenStudy (anonymous):

If 3^-x = 2, what does 3^3x equal?

OpenStudy (anonymous):

you could solve for x using logarithms

OpenStudy (asnaseer):

if:\[3^{-x}=2\implies\frac{1}{3^x}=2\]Now invert both sides - what do you get?

OpenStudy (asnaseer):

invert means takes the reciprocal, e.g. reciprocal of \(\frac{1}{7}\) is \(7\)

OpenStudy (anonymous):

3^x = 2

OpenStudy (asnaseer):

you have only inverted the left-hand-side. you need to invert both sides to keep the equation balanced. e.g. if:\[1+2=3\]then:\[\frac{1}{1+2}=\frac{1}{3}\]and it would be incorrect to state:\[\frac{1}{1+2}=3\]